OpenStudy (anonymous):

[]\*(U^U)*/[]Please help .A number is divided by 3. Then, 5 is added to the quotient, and the result is 8. What is the number.

OpenStudy (imstuck):

\[\frac{ x }{ 3 }+5=8\]

OpenStudy (imstuck):

Multiply all the terms by 3 to get rid of the 3 in that one fraction. Like this:

OpenStudy (imstuck):

\[(3)\frac{ x }{ 3 }+(3)5=(3)8\]That leaves this:

OpenStudy (imstuck):

x+15=24. Now you can easily solve that for x, right?

OpenStudy (anonymous):

\(*>^<)/ thankyou